# Observability-Aware Control for Quadrotor Formation Flight with Range-only Measurement

### v2

@@ 2.4 The Follower Drone Localization Problem with Range-Only Measurements
For a nonlinear system given by Eqs. (2) and (3), state estimators are typically not optimal, and estimation precision can be improved by trajectory optimization. The classical approach is to optimize some metric $\gamma$, such as the trace, determinant, or minimum eigenvalue, of the EKF posterior covariance matrix $\hat{\mathbf{P}}_{k}$ according to [3, 6] as shown:
@@
where $\mathbf{H}_{k}$ is the observation Jacobian and $\check{\mathbf{P}}_{0}$ is the initial prior covariance. However, this approach is not viable for systems where some state variables are not expressed in the observation model. A well-studied example is inertial navigation models augmented by IMU biases [8, 9]. Our CLS is another such example, since the inter-drone velocity $\mathbf{v}$ is not expressed in Eq. (7). In this case, the observation Jacobian contains zero columns at any point of evaluation $\mathbf{x}_{k}$:
$\mathbf{H}_{k}=D\mathbf{h}(\mathbf{x}_{k})=\begin{bmatrix}\mathbf{r}_{k}^{\top}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{1}&\mathbf{0}\end{bmatrix}\in\mathbb{R}^{5\times 10}.$ (10)
@@
This issue arises because $\mathbf{H}_{k}^{\top}\boldsymbol{\Sigma}\mathbf{H}_{k}$ only accounts for observation in timestep $k$, neglecting the evolution of observation along the flow of the system. Recent approaches to approximating the LOG directly address this through Lie derivatives, which motivates a shift towards the LOG for representing localization quality.
@@ 3 Observability Analysis with Range-Only Measurement
In this section, we review the classical nonlinear observability analysis and perform it on our CLS to prove that it satisfies weak local observability. We also introduce a local observability index to prepare for investigating whether weak local observability can be automatically satisfied by observability-aware optimal control. We use the definition from Ref. [18], that a system is observable if and only if the Observability Gramian (OG)
@@
is nonsingular. For a nonlinear system, $\mathbf{W}$ has no closed-form expression and varies with the local point of evaluation, so it is referred to as the LOG [7]. In practice, Hermann and Krener’s [19] rank test are used to detect the observability of a dynamical system. For control-affine systems like Eq. (5), the rth-order nonlinear Observability Matrix is:
@@
The partitions of $\boldsymbol{\mathcal{O}}^{(3)}(\mathbf{x})$ in Eq. (13) are the nonzero rows of $DL\mathbf{h}$, $DL_{\mathbf{f}_{0}}\mathbf{h}$, $DL_{\mathbf{f}_{0}}^{2}\mathbf{h}$, $DL_{\mathbf{f}_{3}\mathbf{f}_{0}}\mathbf{h}$, $DL_{\mathbf{f}_{1}\mathbf{f}_{0}}\mathbf{h}$, and $DL_{\mathbf{f}_{3}}L_{\mathbf{f}_{0}}^{2}\mathbf{h}$, computed by the MATLAB symbolic toolbox [^5] The relevant code can be found in https://github.com/FSC_Lab/observability_aware_controller/symbolic_algebra. By row reduction, it is found to have a rank of 10.
@@
is another observability matrix, parameterized by control inputs but free of the control-affine requirement, as seen in [8].
@@ 4 The Measure of Observability via STLOG
(Ours) Continuous-time observation process $\mathbf{y}(t)$ with GP noise $\boldsymbol{\eta}(t)$ enables inference without relying on discrete observations (Classical) Discrete observations dictate uncertainty propagation(Ours) IC posteriorinferred at any $t\in[0,T]$$\mathbf{x}_{0}$yyttTT
Figure 2: Visualizing initial state estimation by (our) observation process and (classical) discrete observations
@@
Subject to the system model in Eq. (2) and (3), the conditional observation $\mathbf{y}(t)|\mathbf{x}_{0}$ is also a GP, with mean $\mathcal{A}_{t}(\mathbf{x}_{0})$ and covariance $\mathbf{K}(t,t^{\prime})$. This enables us to consider a subproblem of computing the likelihood ratio between two initial conditions, ${p(\mathbf{y}(t)|\mathbf{x}_{0})}/{p(\mathbf{y}(t)|\mathbf{x}_{1})}$. Recall from Mercer’s theorem in Ref. [20] that the covariance kernel $\mathbf{K}(t,t^{\prime})$, under suitable assumptions, defines an integral operator $\mathcal{K}$ in the Hilbert space $L^{2}([0,T],\mathbb{R}^{p})$ of vector-valued square-integrable functions on $[0,T]$. The operator $\mathcal{K}$ has a complete set of orthonormal eigenfunctions $\left\{\mathbf{v}_{n}(t)\right\}_{n=0}^{\infty}$ with respect to the inner product $\langle\cdot,\cdot\rangle$ in $L^{2}([0,T],\mathbb{R}^{p})$}, with corresponding eigenvalues $\left\{\lambda_{n}\right\}_{n=0}^{\infty}$ such that $(\mathcal{K}\mathbf{v}_{n})(t)=\lambda_{n}\mathbf{v}_{n}(t)$.
@@
Next, we introduce a practical noise model, which, when paired with the assumption that the system response $\mathcal{A}:\mathcal{X}\rightarrow C^{\infty}([0,T],\mathbb{R}^{p})$ in Eq. (17) is continuous with respect to the Fréchet topology on $C^{\infty}([0,T],\mathbb{R}^{p})$666 A characterization of the Fréchet topology on $C^{\infty}([0,T],\mathbb{R}^{p})$ is as the metric topology of $d(\mathbf{f}(t),\mathbf{g}(t))=\sum_{k=0}^{\infty}2^{-k}\frac{\rho_{k}(\mathbf{f}-\mathbf{g})}{1+\rho_{k}(\mathbf{f}-\mathbf{g})}$ [21] where $\rho_{k}(\mathbf{f}-\mathbf{g})={\sup}_{t\in[0,T]}\norm{\frac{d^{k}}{dt^{k}}(\mathbf{f}(t)-\mathbf{g}(t))}.$ In particular, all linear systems will have a continuous system response, owing to the existence of an explicit form $\mathcal{A}_{t}(\mathbf{x}_{1})-\mathcal{A}_{t}(\mathbf{x}_{0})=\mathbf{M}(t)\left(\mathbf{x}_{1}-\mathbf{x}_{0}\right)$, for some analytically known smooth matrix-valued function $\mathbf{M}(t)$ on $[0,T]$ [18]. , satisfies the prerequsites of Proposition 4.1. We achieve this by specifying the covariance operator $\mathcal{K}$ spectrally, using the orthonormal Legendre polynomials ${\left\{e_{n}(t)\right\}}_{n=0}^{\infty}$ on $[0,T]$. The detailed rationale is explained in Corollary A.2.1 and Example A.1.
@@
We follow Example A.1 to check that $\mathcal{K}$ in Eq. (21) satisfies the relevant hypotheses in Corollary A.2.1 and apply Proposition 4.1 to express the likelihood ratio as Eq. (19). Then, we further simplify it in Proposition 4.2.
@@
Since $\mathcal{A}(\mathbf{x}_{0})(t)$ is smooth in $t$, its $n$-th Fourer–Legendre coefficient is ${O}(n^{-k})$ for all $k>0$ (see Example A.1). Using the explicit expression for $\mathcal{K}$ in Eq. (21), we compute that, for sufficiently large $N_{c}$,
@@
    The observation horizon $T$ is sufficiently short, but not shorter than the correlation timescale $\delta t$ of the noise $\boldsymbol{\eta}(t)$ in Definition 4.1, so that $\delta t\ll T\ll 1$;777 The STLOG approximation to the LOG is independent of the correlation timescale $\delta t$ per se, but the ordering $\delta t\ll T$ is required for the LOG to emerge from Bayesian state estimation (c.f. Proposition 4.2).
@@
where $L_{\mathbf{f}}=\mathbf{f}\cdot\nabla$ is the Lie derivative on functions [^8] This is consistent with the definition of Lie derivatives in differential geometry (see, e.g. [22]) if we consider $\mathbf{h}$ as a $p$-tuple of scalar functions in the manifold $\mathcal{X}$, whence no contractions involving $D\mathbf{f}$ are necessary. Note, however, that $D\mathbf{h}$ is considered as a $p$-tuple of one-forms on $\mathcal{X}$, and correspondingly $L_{\mathbf{f}}D\mathbf{h}=\mathbf{f}\cdot\nabla(D\mathbf{h})+(D\mathbf{h})(D\mathbf{f})$. The usual identity $L_{\mathbf{f}}D\mathbf{h}=DL_{\mathbf{f}}\mathbf{h}$ holds under this interpretation.. We express $\mathcal{A}_{t}(\mathbf{x}_{0})$ in terms of the derivatives in Eq. (31) taken at $t=0$ to get:
@@
which would be evaluated at $(\mathbf{x}^{*},\mathbf{u};T)$, respectively, the initial condition, control input, and observation time. As such, the STLOG is a novel point-wise representation of observability, in contrast to the traditional LOG, which requires integration over an entire trajectory. It enables local assessment of how the evolution of the system under the current control input contributes to the observability of the current operating state, and aligns with our goal of leveraging observability within predictive control frameworks designed in Section 5.
@@ 5.1 OPC
The goal of observability-aware control is to design a control strategy that refines the estimates of $N$ states $\mathbf{x}(t)$ in any $T_{1},\ldots,T_{k},\ldots,T_{N}$, separated by $\Delta T$ by a discretization scheme. We first propose a law that relates the prior and posterior covariance of state estimation at every $T_{k}$ in a trajectory, by extending Eq. (29) in Proposition 5.1.
@@
where $\boldsymbol{\Phi}_{T_{k},T_{k+1}}=D\left(\boldsymbol{\phi}_{T_{k+1}}\boldsymbol{\phi}^{-1}_{T_{k}}\right)$ is the state transition map from time $T_{k}$ to $T_{k+1}$.999 Similar evolution laws relating the estimation means can also be obtained, but we omit their derivation here since they do not influence the evolution of covariances once the linear approximation around $\mathbf{x}^{*}(t)=\phi_{t}(\mathbf{x}^{*}_{0})$ is admitted, as in Kalman filtering theory for linear systems. Substituting the STLOG approximation in Eq. (34) into Eq. (37) leads to Eq. (35).
Our observability-aware control principle is to collectively maximize the lower bounds to precision improvements, embodied by the minimum eigenvalue of the STLOG in Eq. (36), along a trajectory. This is represented by the objective:
@@
The observability-predictive controller (OPC) optimizes $V$ under a receding horizon framework with the finite horizon taken to be $N=20$ in our software implementation:
@@
where we adopt the shorthands $\mathbf{x}(T_{k})=\mathbf{x}_{k}$, $\mathbf{u}(T_{k})=\mathbf{u}_{k}$. $\mathbf{x}_{cl}$ denotes the current state estimate provided by the CLS, $c$ is a small regularization constant which we take to be $c=10^{-6}$, $d_{k}$ is the inter-drone distance as part of $\mathbf{x}_{k}$, and $d_{\min},d_{\max}$ are the minimum and maximum bounds on $d_{k}$, respectively.

### v3

@@ 2.4 The Follower Drone Localization Problem with Range-Only Measurements
For a nonlinear system such as a quadrotor CLS, state estimators are typically not optimal, and estimation precision can be improved by optimal control. Formulating the optimization problem requires a principled choice of both the precision metric and the system model.
We chose observability, quantified by the LOG and approximated through Lie Derivatives following Ref. [hausman2017observability], as the precision metric because Lie derivatives enable it to reflect how the system configuration, velocity, acceleration, and even higher-order motion lead to informative, evolving observations. Many alternative metrics of positioning quality, such as EKF covariance, the FIM, and the CRLB, rely on first-order linearizations of the observation model. They are generally unsuitable for systems with indirectly observed state components, such as inertial navigation models augmented by biases [hausman2017observability, grebe2021observability] and our CLS, where the inter-drone velocity $\mathbf{v}$ is absent from Eq. (7).
For example, consider the EKF posterior covariance matrix $\hat{\mathbf{P}}_{k}$, used in Refs. [trawny2004optimized, go2024trajectory] and shown below:
@@
$\mathbf{H}_{k}=D\mathbf{h}(\mathbf{x}_{k})=\begin{bmatrix}\mathbf{r}_{k}^{\top}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{1}&\mathbf{0}\end{bmatrix}\in\mathbb{R}^{5\times 10},$ (10)
which causes $\mathbf{H}_{k}^{\top}\boldsymbol{\Sigma}\mathbf{H}_{k}$, also known as the one-step FIM, to have a constantly zero bottom-right block. When the Kalman covariance is updated according to Eq. (9), $\hat{\mathbf{P}}_{k}$ cannot change from $\check{\mathbf{P}}_{0}$ in the velocity subspace, and no change in velocity estimation uncertainty is ever captured. The broader issue, that first-order metrics cannot capture the coupling between motion and observability in indirectly observed states, is shared by other metrics shown in Table 2.
Table 2: Failure modes of first-order metrics under the presence of indirectly observed states.
First-Order Metrics EKF covariance FIM CRLB
[…]
Failure mode Unchanged from prior in subspaces Rank-deficient due to zero blocks Undefined
We chose inter-drone dynamics as our system model to ensure the realizability of optimized motion strategies. The unicycle model, widely used in CLS, is not well-suited to quadrotors, since its motion is coupled to heading, but quadrotor motion is governed by thrust through pitch and roll. Another commonly used simple model is the first-order kinematic system. Such models have proven effective for offline planning of quadrotor trajectories, where spline parameterization [salaris2017online, de2021online] or analytic guidance laws [li2022three] enforce smoothness and dynamical feasibility. However, the present work considers a receding-horizon optimal control formulation, in which control inputs are optimized directly at each timestep. Unaided by trajectory parameterization, kinematic models no longer satisfy implicit smoothness. If observability-based objectives induce sharp variations in the velocity commands to excite the system for informative observations, then quadrotors, which cannot instantaneously change their velocity, may struggle to track them.
We demonstrate this phenomenon in a diagnostic simulation where our leader-follower CLS is instantiated with a first-order kinematic model in a receding-horizon optimal control setting. The follower’s velocity commands at each timestep are optimized under the minimum eigenvalue of the ELOG [krener2009measures] and E2LOG [hausman2017observability] over one predictive horizon, without trajectory parameterization, and bounded below 10 m/s. The accelerations implied by the resulting velocity commands are plotted in Fig. 2. Despite the bounds on velocity, accelerations peaked above $50\mathrm{m/s^{2}}$ and remained above $20\mathrm{m/s^{2}}$ — the maximum achievable by our F450 experimental platforms — for up to 70% and 90% of the time horizon under ELOG and E2LOG, respectively. This is compared to safe acceleration levels below $5\mathrm{m/s^{2}}$, produced by our optimal controller using quadrotor dynamics in subsequent simulations in Sec. 6, and also displayed in Fig. 2.
Figure 2: Translation acceleration required by optimized commands for a first-order kinematic model
These results indicate that, in an optimal control setting without trajectory parameterization, simplified models like first-order kinematic models may yield observability-seeking control commands whose implied accelerations are unattainable by quadrotors. This motivates our choice of the more realistic quadrotor dynamics in Eq. (4) in which thrust and body-rate inputs are optimized directly and are natively compatible with quadrotor low-level controllers.
@@ 3 Observability Analysis with Range-Only Measurement
In this section, we review the classical nonlinear observability analysis and perform it on our CLS to prove that it satisfies weak local observability. We also introduce a local observability index to prepare for investigating whether weak local observability can be automatically satisfied by observability-aware optimal control. We use the definition from Ref. [antsaklis2006linear], that a system is observable if and only if the Observability Gramian (OG)
@@
is nonsingular. For a nonlinear system, $\mathbf{W}$ has no closed-form expression and varies with the local point of evaluation, so it is referred to as the LOG [krener2009measures]. In practice, Hermann and Krener’s [hermann1977nonlinear] rank test is used to detect the observability of a dynamical system. For control-affine systems like Eq. (6), the rth-order nonlinear Observability Matrix is:
@@
The partitions of $\boldsymbol{\mathcal{O}}^{(3)}(\mathbf{x})$ in Eq. (13) are the nonzero rows of $DL\mathbf{h}$, $DL_{\mathbf{f}_{0}}\mathbf{h}$, $DL_{\mathbf{f}_{0}}^{2}\mathbf{h}$, $DL_{\mathbf{f}_{3}\mathbf{f}_{0}}\mathbf{h}$, $DL_{\mathbf{f}_{1}\mathbf{f}_{0}}\mathbf{h}$, and $DL_{\mathbf{f}_{3}}L_{\mathbf{f}_{0}}^{2}\mathbf{h}$, computed by the MATLAB symbolic toolbox [^5] The relevant code can be found in https://github.com/FSC-Lab/observability_aware_controller/symbolic_algebra. By row reduction, it is found to have a rank of 10.
@@
is another observability matrix, parameterized by inputs but free of the control-affine requirement, as seen in [hausman2017observability, yu2015observability].
@@ 4 The Measure of Observability via STLOG
yyttTT (Ours) Continuous-time observation process $\mathbf{y}(t)$ with GP noise $\boldsymbol{\eta}(t)$ enables inference without relying on discrete observations (Classical) Discrete observations dictate uncertainty propagation(Ours) posteriorinferred at any $t\in[0,T]$$\hat{\mathbf{x}}_{0}$
Figure 3: Visualizing initial state estimation by (our) observation process and (classical) discrete observations.
@@
Subject to the system model in Eq. (2) and (3), the conditional observation $\mathbf{y}(t)|\mathbf{x}_{0}$ is also a GP, with mean $\mathcal{A}_{t}(\mathbf{x}_{0})$ and covariance $\mathbf{K}(t,t^{\prime})$. This enables us to consider a subproblem of computing the likelihood ratio between two initial conditions, ${p(\mathbf{y}(t)|\mathbf{x}_{0})}/{p(\mathbf{y}(t)|\mathbf{x}_{1})}$. Recall from Mercer’s theorem in Ref. [seeger2004gaussian] that the covariance kernel $\mathbf{K}(t,t^{\prime})$, under suitable assumptions, defines an integral operator $\mathcal{K}$ in the Hilbert space $L^{2}([0,T],\mathbb{R}^{p})$ of vector-valued square-integrable functions on $[0,T]$. The operator $\mathcal{K}$ has a complete set of orthonormal eigenfunctions $\left\{\mathbf{v}_{n}(t)\right\}_{n=0}^{\infty}$ with respect to the inner product $\langle\cdot,\cdot\rangle$ in $L^{2}([0,T],\mathbb{R}^{p})$}, with corresponding eigenvalues $\left\{\lambda_{n}\right\}_{n=0}^{\infty}$ such that $(\mathcal{K}\mathbf{v}_{n})(t)=\lambda_{n}\mathbf{v}_{n}(t)$.
@@
Next, we introduce a practical noise model, which, when paired with the assumption that the system response $\mathcal{A}:\mathcal{X}\rightarrow C^{\infty}([0,T],\mathbb{R}^{p})$ in Eq. (17) is continuous with respect to the Fréchet topology on $C^{\infty}([0,T],\mathbb{R}^{p})$ [^6] A characterization of the Fréchet topology on $C^{\infty}([0,T],\mathbb{R}^{p})$ is as the metric topology of $d(\mathbf{f}(t),\mathbf{g}(t))=\sum_{k=0}^{\infty}2^{-k}\frac{\rho_{k}(\mathbf{f}-\mathbf{g})}{1+\rho_{k}(\mathbf{f}-\mathbf{g})}$ [reed1980methods] where $\rho_{k}(\mathbf{f}-\mathbf{g})={\sup}_{t\in[0,T]}\norm{\frac{d^{k}}{dt^{k}}(\mathbf{f}(t)-\mathbf{g}(t))}.$ In particular, all linear systems will have a continuous system response, owing to the existence of an explicit form $\mathcal{A}_{t}(\mathbf{x}_{1})-\mathcal{A}_{t}(\mathbf{x}_{0})=\mathbf{M}(t)\left(\mathbf{x}_{1}-\mathbf{x}_{0}\right)$, for some analytically known smooth matrix-valued function $\mathbf{M}(t)$ on $[0,T]$. , satisfies the prerequsites of Proposition 4.1. We achieve this by specifying the covariance operator $\mathcal{K}$ spectrally, using the orthonormal Legendre polynomials ${\left\{e_{n}(t)\right\}}_{n=0}^{\infty}$ on $[0,T]$. The detailed rationale is explained in Corollary A.2.1.
@@
We follow Corollary A.2.1 to check that $\mathcal{K}$ in Eq. (21) satisfies the relevant hypotheses in Corollary A.2.1 and apply Proposition 4.1 to express the likelihood ratio as Eq. (19). Then, we further simplify it in Proposition 4.2.
@@
Since $\mathcal{A}(\mathbf{x}_{0})(t)$ is smooth in $t$, its $n$-th Fourier–Legendre coefficient is ${O}(n^{-k})$ for all $k>0$. Using the explicit expression for $\mathcal{K}$ in Eq. (21), we compute that, for sufficiently large $N_{c}$,
@@
    The observation horizon $T$ is sufficiently short, but not shorter than the correlation timescale $\delta t$ of the noise $\boldsymbol{\eta}(t)$ in Definition 4.1, so that $\delta t\ll T\ll 1$; [^7] The STLOG approximation to the LOG is independent of the correlation timescale $\delta t$ per se, but the ordering $\delta t\ll T$ is required for the LOG to emerge from Bayesian state estimation (c.f. Proposition 4.2).
@@
where $L_{\mathbf{f}}=\mathbf{f}\cdot\nabla$ is the Lie derivative on functions [^8] This is consistent with the definition of Lie derivatives in differential geometry if we consider $\mathbf{h}$ as a $p$-tuple of scalar functions in the manifold $\mathcal{X}$, whence no contractions involving $D\mathbf{f}$ are necessary. Note, however, that $D\mathbf{h}$ is considered as a $p$-tuple of one-forms on $\mathcal{X}$, and correspondingly $L_{\mathbf{f}}D\mathbf{h}=\mathbf{f}\cdot\nabla(D\mathbf{h})+(D\mathbf{h})(D\mathbf{f})$. The usual identity $L_{\mathbf{f}}D\mathbf{h}=DL_{\mathbf{f}}\mathbf{h}$ holds under this interpretation.. We express $\mathcal{A}_{t}(\mathbf{x}_{0})$ in terms of the derivatives in Eq. (31) taken at $t=0$ to get:
@@
which would be evaluated at $(\mathbf{x}^{*},\mathbf{u};T)$, respectively, the initial condition, control input, and observation time. As such, the STLOG is a novel point-wise representation of observability, which enables local assessment of how the evolution of the system under the current control input contributes to the observability of the current operating state, and aligns with our goal of leveraging observability within predictive control frameworks designed in Section 5.
@@ 5.1 OPC
The goal of observability-aware control is to design a control strategy that refines the estimates of $N$ states $\mathbf{x}(t)$ in any $T_{1},\ldots,T_{k},\ldots,T_{N}$, separated by $\Delta T$ in a discretization scheme. We first propose a law that relates the prior and posterior covariance of state estimation at every $T_{k}$ in a trajectory, by extending Eq. (29) in Proposition 5.1.
@@
where $\boldsymbol{\Phi}_{T_{k},T_{k+1}}=D\left(\boldsymbol{\phi}_{T_{k+1}}\boldsymbol{\phi}^{-1}_{T_{k}}\right)$ is the state transition map from time $T_{k}$ to $T_{k+1}$. [^9] Similar evolution laws relating the estimation means can also be obtained, but we omit their derivation here since they do not influence the evolution of covariances once the linear approximation around $\mathbf{x}^{*}(t)=\phi_{t}(\mathbf{x}^{*}_{0})$ is admitted, as in Kalman filtering theory for linear systems. Substituting the STLOG approximation in Eq. (34) into Eq. (37) leads to Eq. (35).
Our observability-aware control principle is to collectively maximize the lower bounds to precision improvements, embodied by the minimum eigenvalue of the STLOG in Eq. (36), along a trajectory. This corresponds to a safety-oriented, robustness-driven design: by maximizing the smallest eigenvalue, the controller prioritizes the least observable state direction, guaranteeing a minimum reduction in estimation uncertainty and preventing severe loss of observability that could lead to CLS divergence. This is represented by the objective:
@@
The observability-predictive controller (OPC) optimizes $V$ under a receding horizon framework with a horizon $N$:
@@
where $\mathbf{x}_{0}=\mathbf{x}_{cl}$ is the feedback of the current state, consisting of the inter-drone position, orientation, and velocity, estimated by a CLS. $\mathbf{x}(T_{k})=\mathbf{x}_{k}$ and $\mathbf{u}(T_{k})=\mathbf{u}_{k}$ are the states and input at time $T_{k}$ along the trajectory to be optimized. Following the receding-horizon paradigm, only the first control input $\mathbf{u}_{0}$, consisting of thrust and body rates, is applied to the system. Lastly, $c$ is a small regularization constant which we take to be $c=10^{-6}$.
Unlike classical formation control, Eq. (39) does not penalize deviations from a specific formation or predefined positions between the leader and follower. Instead, at every step $k$, the leader-follower distance $d_{k}$ is restricted to lie within the bounds $d_{\min},d_{\max}]$ to maintain ranging measurements and connectivity. In practice, the OPC operates as a high-level receding-horizon controller that receives the current CLS state estimate $\mathbf{x}_{cl}$ as feedback and solves for optimal thrust and body-rate commands included in $\mathbf{u}_{k}$, which are tracked by quadrotor low-level attitude controllers.